Suppose that the time required to complete a 1040R tax form is normally distributed with a mean of 100 minutes and a standard deviation of 20 minutes. What proportion of 1040R tax forms will be completed in more than 127 minutes? Round your answer to at least four decimal places.


1
Expert's answer
2021-05-25T15:58:00-0400

"\mu=100,\sigma=20"


Proportion of 1040R tax forms will be completed in more than 127 minutes-


"P(X>127)=P(z>\dfrac{127-100}{20})=P(z>\dfrac{27}{20})=P(z>1.35)=0.0885"


Hence The proportion is 8.85%.